OpenStudy (anonymous):

Hexagon A is a regular hexagon. The total length of all the sides of the hexagon is 36 inches. Hexagon A is dilated about its center to create Hexagon B. The length of each side of Hexagon B is 19 1/2 inches. By what factor was Hexagon A dilated to create Hexagon B?

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } or \frac{ 4 }{ 13 } or 2\frac{ 1 }{ 4 } or 3\frac{ 1 }{ 4 }\]

OpenStudy (cwrw238):

divide 19 1/2 by 6 ( 6 because all 6 sides = 36 so one side = 6)

OpenStudy (cwrw238):

can you do this division Tati_Li?